OpenStudy (kate1624):

Could somebody help me out I don't trust my answers. 1) -2/3x-9<9/10 My answer: x< -99/10 2) 4x+6<-6 My answer: x<-3 3) q+12-2(q-22)>0 My answer: Idek

jimthompson5910 (jim_thompson5910):

First one is \[\Large -\frac{2}{3}x-9<\frac{9}{10}\] right?

OpenStudy (kate1624):

yes

jimthompson5910 (jim_thompson5910):

the answer you got is incorrect

OpenStudy (kate1624):

figures

jimthompson5910 (jim_thompson5910):

however, the answer you got for 4x+6<-6 is correct

jimthompson5910 (jim_thompson5910):

to make -2/3x-9<9/10 simpler, you can multiply both sides by the LCD 30 -2/3x-9<9/10 30(-2/3x-9)<30*(9/10 ) 30(-2/3x) + 30(-9) < 27 -20x - 270 < 27 hopefully you see how to finish up to solve for x

OpenStudy (kate1624):

o shoot no it was actually supposed to be -2/5 instead of -2/3 I was looking at the wrong problem and added them together instead :/

jimthompson5910 (jim_thompson5910):

so, \[\Large -\frac{2}{5}x-9<\frac{9}{10}\] ???

OpenStudy (kate1624):

But yes I do understand how you got that answer

OpenStudy (kate1624):

yes

jimthompson5910 (jim_thompson5910):

x< -99/10 is close, but still not correct

jimthompson5910 (jim_thompson5910):

-2/5x-9<9/10 10*(-2/5x-9)<10*(9/10) ... Multiply both sides by the LCD 10. 10*(-2/5x) + 10*(-9) <10*(9/10) -4x - 90 < 9 solve for x

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (kate1624):

-99/4 or -24.75

jimthompson5910 (jim_thompson5910):

x > -99/4 OR x < -99/4 ??

OpenStudy (kate1624):

x > -99/4

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

# 3 q+12-2(q-22) > 0 q+12-2q-2(-22) > 0 ... distribute q + 12 - 2q + 44 > 0 -q + 56 > 0 I'll let you finish up